# Learning Keypoints for Robotic Cloth Manipulation using Synthetic Data

### v1

@@ VI-A Main Results
Figure 3: Illustration of failure modes of the detectors. From left to right, each row shows ground truth keypoints and the output of detectors trained on real data, synthetic data and on both. The first row illustrates how the real baseline produces incomplete and inconsistent keypoints, whereas the second row shows how the detectors still struggle with folds. In the third row, the detectors are confused by an open zipper and mistake it for a leg of the shorts.
@@
mAP2,4,8 (normal-$\uparrow$) mAP2 (normal-$\uparrow$)
@@
T-shirt 58.2 81.8 43.8 69.5
Towel 83.2 84.7 77.7 78.5
Shorts 51.4 79.8 44.0 71.4
All 64.3 82.1 55.2 73.1
@@
The performance of all models, as measured on the aRTF test splits, is given in Table II. Over all cloth types, the baseline model that is only trained on real data obtains an mAP2,4,8 of $59.859.8$ and an mAP2 of $44.644.6$. Training on synthetic data outperforms this baseline by 4 and 9 percentage points (pp ) respectively. Fine-tuning on the aRTF data results in an mAP2,4,8 of $74.274.2$ and an mAP2 of $63.763.7$, which is an improvement of 10 pp compared to training on synthetic data only. The performance on T-shirts and shorts is rather similar, whereas the results for towels are respectively 10 and 20 pp above average for both metrics. A few examples of the output of the models trained on a combination of the aRTF train split and synthetic data can be found in Fig. 1. Using these keypoints, folding policies can be scripted [13, 5].
The above results clearly show the added value of the synthetic data. At the same time, the results in Table II also indicate that the keypoints are still not always detected correctly, as this would result in an mAP close to 100%. To provide more insight into the failures of the detectors, we show predictions of the models trained on each data source on a number of samples from the aRTF test split in Fig. 3. Each example illustrates a common failure mode: The first row illustrates how training only on the aRTF train split results in models that confuse semantic locations and return many false positives, which is also visible in the second example. The keypoint detectors still struggle with folds, as can be seen in the second example. Despite many efforts to tailor the distribution of mesh deformations in the synthetic data (see Section IV-B), we have not managed to completely solve this problem. The final example in Fig. 3 illustrates how all models have difficulties with a particular configuration of a short in which the opened zipper is confused for a leg. Similar confusions occur for T-shirts as well.
The lack of further improvement when adding more synthetic data and the performance improvement when fine-tuning on the aRTF data already suggest that there is a reality gap for the synthetic data. To further quantify this gap, we have evaluated both the sim-to-sim performance and the sim-to-real performance of the models trained on synthetic data. These results can be found in table III. Across all categories, the gap is about 20pp , which confirms there is a substantial reality gap. Interestingly, for towels the sim and real performance are very close, suggesting that the gap is smaller for this category. This also corresponds with the results from Table II, where the performance increase associated with fine-tuning on the aRTF data was lower than for the other categories.
@@ VI-B Comparing procedures for obtaining cloth meshes
In this experiment, we motivate the use of single-layer meshes for synthetic data generation by comparing three different procedures to generate the cloth meshes:
@@
    Inspired by [5] we have taken 5 meshes from the Cloth3D dataset [25]. We have triangulated the meshes and cleaned up their UV maps. Using Nvidia Flex, we generated a flattened version of these meshes, by dropping them from a certain height. We then manually labeled all keypoints by selecting the corresponding vertices on the mesh. Similar to [14], we have created 10 variants of each mesh by rescaling it, leading to a total of 50 different meshes. Using these meshes, we generate deformed configurations with the procedure discussed in Section IV-B.
@@
    We procedurally generate single-layer meshes using a 2D template as described in Section IV-A1. We then create deformations using the same procedure as above. As these meshes and their annotations are generated procedurally, we can generate arbitrary amounts of meshes, unlike for the cloth3d procedure.
@@
Figure 4: Examples of the different cloth mesh procedures that were considered. Left to right: undeformed single-layer mesh, cloth3d mesh, single-layer mesh. Though less realistic, using the single-layer meshes results in the best performance.
@@
Cloth Mesh Procedure mAP2,4,8 (normal-$\uparrow$) mAP2 (normal-$\uparrow$)
Cloth3D subset 43.4 29.1
single-layer 54.3 38.7
single-layer undeformed 53.8 39.0
@@ VI-C Comparing different cloth materials
In this experiment, we compare several different cloth material configurations, to motivate the use of random materials in Section 2:
@@
An example of each material can be found in Fig. 5. We have generated datasets of 5,000 images for these materials, using the procedure from Section IV-C. The performance of the keypoint detectors trained on these datasets can be found in table V. We observe that uniform color materials result in the lowest performance. This is not surprising as the aRTF dataset contains many non-uniform colored cloth items. It is also consistent with findings in our previous work [19], where we also observed a decreased performance for non-uniform cloth pieces. Perhaps more surprisingly, the random materials perform better than the tailored materials even though the latter look more plausible.
@@
Cloth Materials mAP2,4,8 (normal-$\uparrow$) mAP2 (normal-$\uparrow$)
PolyHaven random 54.3 38.7
cloth-tailored 49.0 36.5
uniform color 35.1 25.5
@@ VII Discussion
Not only do the assets need to be improved, particle-based simulators such as Nvidia Flex [16] have fundamental limitations. Next to the limited realism of the cloth-physics [3, 40], we have also found the inherent intertwining of mesh resolution with cloth thickness and deformation granularity hindering. Simulators such as C-IPC [41] provide improved realism [13], though at the cost of additional computational complexity.
@@
Next to the sim-to-real gap, we have observed that the keypoint detectors still struggle with folds and severe deformations. This results in predictions that are between themselves not consistent and false positives on the apparent cloth edges. It is very well possible that these issues can be resolved with improved data generation, but we think that the models would benefit from more geometric reasoning capabilities. On the other hand, we are not convinced that forcing a model to predict keypoints from a static image of heavily deformed clothes lying on a surface is the best way forward: this might be an artificially hard task. Using more interactive perception as in [12] would allow the system to collect more information. Improvements to the generalization of unfolding systems could also reduce the need to handle deformed pieces in the first place.
@@ VIII Conclusion
In this work, we have described a pipeline to generate synthetic images of clothes to advance the generalization of robotic cloth manipulation. By training a keypoint detector on images generated by this pipeline we have validated its usefulness. The resulting models and the aRTF dataset can be used to enable various tasks such as folding or ironing. Furthermore, the pipeline can be easily adapted to train models for other tasks or cloth types.

### v2

@@ VI-A Main Results
mAP2,4,8 ($\uparrow$) AKD($\downarrow$)
@@
T-shirt 58.2 81.8 14.0 5.8
Towel 83.2 84.7 13.1 7.4
Shorts 51.4 79.8 27.3 7.4
All 64.3 82.1 18.1 6.8
@@
The performance of all models, as measured on the aRTF test splits, is given in Table II. Over all cloth types, the baseline model that is only trained on real data obtains an mAP2,4,8 of $59.859.8$ and an AKD of $16.816.8$ pixels. Training on synthetic data results in an improved mAP2,4,8 ($64.364.3$) but a lower AKD of $18.118.1$ pixels, caused by a longer tail of the keypoint distances for the model trained on synthetic data. Fine-tuning on the aRTF data results in an mAP2,4,8 of $74.274.2$ and an AKD of only $8.78.7$ pixels, almost halving the average distance compared to training on real or synthetic data only. When comparing between categories the performance is best for the towels on both metrics. A few examples of the output of the models trained on a combination of the aRTF train split and synthetic data can be found in Fig. 1. We have also added a video to the repository where we visualize the keypoint predictions while interacting with various clothes. Using the keypoint detectors, folding policies can be scripted [13, 5]. We demonstrate this by folding T-shirts using scripted motions based on the detected keypoints. We do not report quantitative results, for which our experimerent was not extensive enough, but have included a video in the project repository.
The above results clearly show the added value of the synthetic data. At the same time, Table II shows that the detectors are not perfect. To provide more insight into the failures of the detectors, we show the predicted keypoints of the models trained on each data source on a number of samples from the aRTF test split in Fig. 3. Each example illustrates a common failure mode: The first row illustrates how training only on the aRTF train split results in models that confuse semantic locations and return many false positives, which is also visible in the second example. The keypoint detectors still struggle with folds, as can be seen in the second example. Despite many efforts to tailor the distribution of mesh deformations in the synthetic data (see Section IV-B), we have not managed to completely solve this problem. The final example in Fig. 3 illustrates how all models have difficulties with a particular configuration of a short in which the opened zipper is confused for a leg. Similar confusions occur for T-shirts as well.
The lack of further improvement when adding more synthetic data and the performance improvement when fine-tuning on the aRTF data already suggest that there is a reality gap for the synthetic data. To measure this, we have evaluated both the sim-to-sim performance and the sim-to-real performance of the models trained on synthetic data. These results can be found in table III. Across all categories, the difference is about 20pp for the mAP and 11 pixels for the AKD, which confirms there is still a substantial reality gap.
@@ VI-B Comparing procedures for obtaining cloth meshes
In this experiment, we motivate the use of single-layer meshes for synthetic data generation by comparing three different procedures to generate T-shirt meshes:
@@
    Inspired by [5] we have taken 5 T-shirt meshes from the Cloth3D dataset [25]. We have triangulated the meshes and cleaned up their UV maps. Using Nvidia Flex, we generated a flattened version of these meshes, by dropping them from a certain height. We then manually labeled all keypoints by selecting the corresponding vertices on the mesh. Similar to [14], we have created 10 variants of each mesh by rescaling it, leading to a total of 50 different meshes. Using these meshes, we generate deformed configurations with the procedure discussed in Section IV-B.
@@
    We procedurally generate single-layer T-shirt meshes using a 2D template as described in Section IV-A1. We then create deformations using the same procedure as above. As these meshes and their annotations are generated procedurally, we can generate arbitrary amounts of meshes, unlike for the cloth3d procedure.
@@
Figure 4: Examples of the considered cloth mesh procedures. Left to right: undeformed single-layer mesh, Cloth3d mesh, single-layer mesh. Though less realistic, using the single-layer meshes results in the best performance.
@@
Cloth Mesh Procedure mAP2,4,8 ($\uparrow$) AKD($\downarrow$)
Cloth3D subset 43.4 20.6
single-layer 54.3 15.3
single-layer undeformed 53.8 17.9
@@ VI-C Comparing different cloth materials
In this experiment, we compare different cloth material configurations for rendering:
@@
An example of each material can be found in Fig. 5. We have generated datasets of 5,000 images for these materials, using the procedure from Section IV-C. The performance of the keypoint detectors trained on these datasets can be found in table V. We observe that uniform color materials result in the lowest performance. This is not surprising as the aRTF dataset contains many non-uniform colored cloth items. It is also consistent with findings in our previous work [19], where we also observed a decreased performance for non-uniform cloth pieces. Perhaps more surprisingly, the random materials perform better than the tailored materials even though the latter look more plausible. We therefore use random materials for all other experiments in this paper.
@@
Cloth Materials mAP2,4,8 ($\uparrow$) AKD($\downarrow$)
PolyHaven random 54.3 15.3
cloth-tailored 49.0 22.1
uniform color 35.1 35.4
@@ VII Discussion
Not only do the assets need to be improved, particle-based simulators such as Nvidia Flex [16] have fundamental limitations. Next to the limited realism of the cloth physics [3, 40], we have also found the inherent intertwining of mesh resolution with cloth thickness and deformation granularity hindering. Simulators such as C-IPC [41] provide improved realism [13], though at the cost of additional computational complexity.
@@
Though we firmly believe in the benefits of synthetic data and their use for robotic manipulation, we are not convinced that forcing a model to predict keypoints from a static image of heavily deformed clothes lying on a surface is the best way forward: this might be an artificially hard task. Using more interactive perception as in [12] would allow the system to collect more information. Finally, improvements to the generalization of unfolding systems could also reduce the need to handle deformed pieces in the first place.
@@ VIII Conclusion
In this work, we have described a pipeline to generate synthetic images of clothes with the goal of improving generalization of robotic cloth manipulation systems. By training keypoint detectors on images generated by this pipeline we have validated its usefulness. The resulting models and the aRTF dataset can be used to enable various tasks such as folding or ironing. Furthermore, the pipeline can be easily adapted to train models for other tasks or cloth types.